# CLST: Cold-Start Mitigation in Knowledge Tracing by Aligning a Generative Language Model as a Students’ Knowledge Tracer

### v1

@@ 1 Introduction
In this study, we propose a framework called “cold-start mitigation in KT by a aligning generative language model as a students’ knowledge tracer (CLST).” We conducted experiments to determine whether a generative LLM aligned with KT could perform well in cold-start scenarios. Initially, we collected data for a variety of subjects, including mathematics, social studies, and science. The KT task was then framed as a natural language processing (NLP) task by expressing problem-solving data in natural language, and the formatted KT dataset was used to fine-tune the generative LLM model. Finally, multiple experiments were conducted to investigate CLST’s performance in cold-start situations. This study was conducted to address the following research questions (RQs):
@@ 4.2.2 Effectiveness of fine-tuning on model’s reliability
The x-axis of each heat map represents the interaction (KC id, correctness) at each time step, while the y-axis represents the KC of the questions that the student primarily answered. In other words, the value of each cell indicates the change in the mastery level for the corresponding KC after the interaction occurred. The x-axis of the line plot (lower part of Figures 5 to 8) likewise represents the interaction (KC id, correctness) at each time step, with the y-axis representing mastery level values; thus, each line represents the mastery level of the corresponding KC.
@@
Similarly, in the learning trajectory of another math student shown in Figure 6, the mastery levels of related KCs, namely KC-25 (division of fractions) and KC-92 (addition and subtraction of fractions), as well as KC-20 (percent of) and KC-37 (conversion of fractions, decimals, and percents),
@@
Ultimately, this experiment reveals the following answer to RQ4: Models trained with the CLST are more generalizable across domains; therefore, when building an ITS in situations where data is scarce and sufficient student–exercise interactions cannot be obtained, the CLST can be a highly promising choice.
@@ 5 Conclusion
The results of this study have practical implications, serving as a guide for educational institutions and EdTech companies interested in fostering personalized learning through the development of ITS and data collection via KT tasks. Additionally, it offers theoretical implications by shedding light on strategies for leveraging generative LLMs in the KT domain.

### v2

@@ 1 Introduction
In this study, we propose a framework called “cold-start mitigation in KT by aligning a generative language model as a students’ knowledge tracer (CLST).” We conducted experiments to determine whether a generative LLM aligned with KT could perform well in cold-start scenarios. Initially, we collected data for a variety of subjects, including mathematics, social studies, and science. The KT task was then framed as a natural language processing (NLP) task by expressing problem-solving data in natural language, and the formatted KT dataset was used to fine-tune the generative LLM. Finally, multiple experiments were conducted to investigate CLST’s performance in cold-start situations. This study was conducted to address the following research questions (RQs):
@@ 4.2.2 Effectiveness of fine-tuning on model’s reliability
The x-axis of each heat map represents the interaction (KC id, correctness) at each time step, while the y-axis represents the KC of the questions that the student primarily answered. In other words, the value of each cell indicates the change in the mastery level for the corresponding KC after the interaction occurred. The x-axis of the line plot (lower part of Figures) likewise represents the interaction (KC id, correctness) at each time step, with the y-axis representing mastery level values; thus, each line represents the mastery level of the corresponding KC.
@@
Similarly, in the learning trajectory of another math student shown in Figure 6, the mastery levels of related KCs, namely KC-25 (division of fractions) and KC-92 (addition and subtraction of fractions), as well as KC-20 (percent of) and KC-37 (conversion of fractions, decimals, and percents), displayed similar trends.
@@
Ultimately, this experiment reveals the following answer to RQ4: Models trained with the CLST are more generalizable across domains; therefore, when building an ITS in situations where data is scarce and sufficient student–exercise interactions cannot be obtained, the CLST can be a promising choice.
@@ 5 Conclusion
The results of this study have practical implications, serving as a guide for educational institutions and EdTech companies interested in fostering personalized learning through the development of ITS. Additionally, it offers theoretical implications by shedding light on strategies for leveraging generative LLMs in the KT domain.